MODS1 Commissioning Run 5: UTC 2011 February 21
Observer(s): R. Pogge, J. Antognini, & P. Martini (OSU), M. Wagner (LBTO)
Telescope Operator: D. Gonzalez-Huerta (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Urban (LBTO)
Instrument Support: J. Morris (LBTO)
No observing tonight after a day of blowing freezing fog. High humidity all night and rime ice covering every surface outside. No chance of
Doing some work on examining data from previous nights and tweaking up and testing scripting and mask alignment code. Some more tests of
The dewar hold-time test continues from yesterday afternoon at 1600MST. We won't fill the dewars again until we see indications from the
temperature and vacuum gauges that the dewars are warming up. No issue for observing prep since we are very likely not going to opening
tonight because of bad weather and ice accumulation on everything.
All times are UTC
unless otherwise noted.
- There is little snow but a generous and very beautiful frosting of rime ice on all the trees and buildings on the LBT summit. Humidity is 100% with a temperature of -8°C. Still very windy with speeds of 15-20m/s out of the S. The anemometer is giving occasional screwy readings due to rime ice build up overwhelming its heater.
- Seeing the moon through fog, humidity is still 100% and the wind is 12-15m/s out of the SSW with gusts up to 19m/s. Getting colder, now -11°C, falling at a rate of about -1°C per hour. Everything is still iced up, even the all-sky camera at the VATT, but the anemometer seems to have cleared itself of ice and is better behaved than before.
Dewar Hold-Time Experiment
The MODS CCD dewars were last filled on MST Feb 20 1800 (UTC Feb 19 2300). Red started warming up at 0500 UTC Feb 21, and Blue at 0800 UTC.
The approximate dewar hold-times are
MODS1 Red CCD: 30 hours
MODS1 Blue CCD: 33 hours
- 20 Feb 2011